OpenStudy (anonymous):

Which expression can be used to find the slope of a line containing the points (–3, 2) and (7, –1)? mc003-4.jpg

jimthompson5910 (jim_thompson5910):

You will use the slope formula \[\Large m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

jimthompson5910 (jim_thompson5910):

(x1,y1) is your first point (x2,y2) is the second point

OpenStudy (anonymous):



OpenStudy (anonymous):

no dear you need to subtract, not put commas there

OpenStudy (anonymous):

\[m=\frac{2-(-1)}{-3-7}\]

jimthompson5910 (jim_thompson5910):

In this case, y1 = 2 y2 = -1 when you subtract, you get y2 - y1 = -1-2 = -3 do the same for the x's (same order) and divide. Reduce as much as possible.

OpenStudy (anonymous):

Can you just give me the answer ?

jimthompson5910 (jim_thompson5910):

what do you get when you subtract the x values?

OpenStudy (anonymous):

the question asks for "which expression" not the answer do you have choices for expressions?